A conference hall, the number of seats in each row is 10 fewer than the number of rows. The number of seats in the hall are 600. Find the number of seats in each row.

Grade:11

1 Answers

Bipin Jung Thakuri
5 years ago
let no of seats be x and rows be y
here,
x=y-10
y=x+10
total no of seats =xy =x(x+10)
x(x+10)=600
on solving it we get, x=20